### v2

# Degrees of interior polynomials and parking function enumerators

Abstract.
The interior polynomial of a directed graph is defined as the $h^{*}$-polynomial of the graph’s (extended) root polytope, and it displays several attractive properties. Here we express its degree in terms of the minimum cardinality of a directed join, and give a formula for the leading coefficient. We present natural generalizations of these results to oriented regular matroids; in the process we also give a facet description for the extended root polytope of a regular oriented matroid.
By duality, our expression for the degree of the interior polynomial implies a formula for the degree of the parking function enumerator of an Eulerian directed graph (which is equivalent to the greedoid polynomial of the corresponding branching greedoid). We extend that result to obtain the degree of the parking function enumerator of an arbitrary rooted directed graph in terms of the minimum cardinality of a certain type of feedback arc set.
1. Introduction
In this paper we compute the degrees of two interrelated (in fact, in an appropriate sense, dual) graph and matroid polynomials. In particular, we show that they can be expressed using common graph/matroid theoretic concepts.
1.1. Interior polynomials
The first type of polynomial we deal with is the interior polynomial of a directed graph. This notion generalizes some well-studied graph invariants, most notably the specialization $T(x,1)$ of the Tutte polynomial, as well as its extension to hypergraphs [10]. It is defined as the $h^{*}$-vector of the (extended) root polytope
[…]
associated to the digraph $G=(V,E)$. Please see Section 2 for detailed definitions. In this paper we show that the degree and leading coefficient of the interior polynomial have meaningful connections to the graph structure.
In a directed graph, we call a cut directed if all of its edges point toward the same shore. An edge set in a digraph is called a directed join, or dijoin for short, if it intersects every directed cut. Our first main result is the following.
Theorem 1.1.
Let $G=(V,E)$ be a connected digraph. Then the degree of the interior polynomial of $G$ is equal to $|V|-1-\nu(G)$, where $\nu(G)=\min\{|K|\mid K\subseteq E\text{ is a dijoin of }G\}$.
Recall that by a theorem of Lucchesi and Younger [16], the quantity $\nu(G)$ above is also the maximal number of edge-disjoint directed cuts in $G$. Furthermore, if the underlying undirected graph of $G$ is 2-edge connected, then $\nu(G)$ is the minimal number of edges whose reversal yields a strongly connected orientation of $G$ [8, Proposition 9.7.1].
We also express the leading coefficient of the interior polynomial $I_{G}$ of $G$.
Definition 1.2.
Let $G$ be a directed graph, and $F$ a subset of its edges. We call the vector $\mathbf{z}=\sum_{v\in V}z_{v}\mathbf{1}_{v}\in\mathbb{Z}^{V}$ with
$z_{v}=|\{e\in F\mid v\text{ is the head of }e\}|-|\{e\in F\mid v\text{ is the tail of }e\}|$
the net degree vector of $F$.
Theorem 1.3.
The leading coefficient of $I_{G}$ equals the number of vectors that can be obtained as the net degree vector of a minimal cardinality dijoin in $G$.
Let us remark that the constant and linear terms of interior polynomials can also be expressed by using basic Ehrhart theory. First of all the constant term of any $h^{*}$-polynomial is $1$. Furthermore, in all cases when $G$ is connected, the dimension of the extended root polytope is $|V|-1$. This is because, on the one hand, $\dim{\tilde{\mathcal{Q}}}_{G}\leq|V|-1$ is obvious from the definition; on the other hand, any spanning tree $T$ of $G$ gives rise to a simplex ${\tilde{\mathcal{Q}}}_{T}$ of dimension $|V|-1$ inside ${\tilde{\mathcal{Q}}}_{G}$. In fact, the extended root polytope can be triangulated by such simplices, which also happen to be unimodular. Hence $I_{G}$ agrees with the $h$-vector of the triangulation, and by the relation between $f$- and $h$-vectors, the coefficient of the linear term in $I_{G}$ is [^1] Here $|E|+1$ is the number of generators of ${\tilde{\mathcal{Q}}}_{G}$, all of whom become vertices in the unimodular triangulation, and $|V|-1$ is the dimension. $|E|+1-(|V|-1)-1=|E|-|V|+1$, that is the genus (also known as the nullity, cyclomatic number, or simply the first Betti number) of $G$.
Figure 1. A directed graph with five disjoint directed cuts.
Example 1.4.
Figure 1 shows a directed graph $G$ whose interior polynomial is $I_{G}(x)=4x^{2}+3x+1$. This can be checked directly in various ways, one of which is mentioned in the next Remark. Note that the coefficient of $x$ indeed agrees with $|E|-|V|+1=3$. As to the leading coefficient, first notice that the edge set of $G$ in this case is the disjoint union of five directed cuts, whence $\nu(G)\geq 5$. In fact $\nu(G)=5$ since it is easy to find dijoins of cardinality $5$ (actually, there are $18$ of them). Thus Theorem 1.1 confirms that $\deg I_{G}=8-1-5=2$ and by Theorem 1.3, the leading coefficient should indeed be $4$ because the minimal dijoins induce the following four net degree vectors:
111111-1- 1222211-1- 111-1- 1111122-2- 21111
; 222211-1- 1111111-1- 111-1- 1111122-2- 21111 ; 111111-1- 1222211-1- 122-2- 2111111-1- 11111 ; 222211-1- 1111111-1- 122-2- 2111111-1- $1$
Remark 1.5.
Let us explain a simple case that inspired Theorem 1.1. The interior polynomial was first introduced by Kálmán [10], for hypergraphs. By the main result of [11], that definition corresponds to the situation when $G=(U,W;E)$ is a bipartite graph with classes $U$ and $W$, and each edge is oriented toward $W$. (The graph $G$ of Example 1.4 is of this type with $|U|=|W|=4$, $\nu(G)=5$, and $\deg I_{G}=2$. In particular, its interior polynomial may be evaluated by using [10, Definition 5.3].)
The hypergraphical approach to the interior polynomial $I_{G}$ easily yields (see [10, Proposition 6.1]) that (for graphs $G$ as above)
(1.1) $\deg I_{G}\leq\min\{|U|,|W|\}-1.$
By considering star-cuts, it is also obvious that
(1.2) $\nu(G)\geq\max\{|U|,|W|\}.$
In this special case, Theorem 1.1 is the statement that the defects in the inequalities (1.1) and (1.2) coincide. This realization, made jointly with András Frank, predates Theorem 1.1 and was subsumed by it later. Regrettably, neither (1.1) nor (1.2) seems to have a generalization to wider classes of digraphs.
We note that interior polynomials of hypergraphs have another, rather different generalization, this time to polymatroids instead of directed graphs (see [10], as well as [4] for a two-variable version). As far as we know, the connection to $h^{*}$-polynomials does not extend to the polymatroid setting, and a degree formula for the interior polynomial of a polymatroid remains unknown. We do not touch on polymatroids in this paper.
By analogy to directed graphs, one can introduce the (extended) root polytope of a regular oriented matroid and define the interior polynomial as its $h^{*}$-vector. (For details, see Section 4.) It turns out that all arguments readily generalize to this case, and one is able to obtain analogous statements regarding the degree and leading coefficient.
Theorem 1.6.
Let $M$ be a regular oriented matroid of rank $r$. Then the degree of the interior polynomial $I_{M}$ of $M$ is $r-\nu(M)$, where $\nu(M)=\min\{|K|\mid K\text{ is a dijoin of }M\}$.
Theorem 1.7.
Take an arbitrary totally unimodular representing matrix $A=[\mathbf{a}_{i}]_{i\in E}$ for the regular oriented matroid $M$. The leading coefficient of $I_{M}$ equals the number of vectors that can be obtained as $\sum_{i\in K}\mathbf{a}_{i}$, where $K$ is a minimal cardinality dijoin of $M$.
Even though Theorems 1.6 and 1.7 imply Theorems 1.1 and 1.3, respectively, we include separate proofs for the benefit of those readers who are less interested in oriented matroids. A facet description of the extended root polytope is provided in Proposition 4.1. We remark that D’Ali, Juhnke-Kubitzke and Koch [7] found a similar facet description in a special case, namely the one that corresponds to the symmetric edge polytope of a graph.
It is interesting to compare interior polynomials for different orientations of a given undirected graph. In an earlier paper [12] we highlighted a special case of this problem. Namely, we conjectured that for a bipartite graph, among its so-called semi-balanced orientations (see Section 2), the one in which all edges point toward the same class has a coefficientwise minimal interior polynomial. (There are of course two such orientations but their extended root polytopes are isometric and thus their interior polynomials identical.) Semi-balanced orientations are relevant because their root polytopes are the facets of the symmetric edge polytope of the bipartite graph. Here we prove a weakened version of this conjecture, addressing only the degree: We show that, in fact among all orientations of a bipartite graph, the one with each edge pointing toward one of the classes has an interior polynomial of minimal degree. The analogous problem remains open for non-bipartite graphs and for regular matroids, where we do not even have a candidate for the degree-minimizing orientation. We do however conjecture that in these broader cases, too, there is an orientation that coefficientwise minimizes the interior polynomial.
1.2. Greedoid polynomials
The other main object of this paper is the greedoid polynomial, introduced by Björner, Korte, and Lovász [5]. It is set up in such a way that interesting information is contained in its lowest-degree term. (The degree of the leading term depends only weakly on the greedoid, namely it is the size of the ground set minus the rank. For exact definitions, see Section 5.)
In the special case of branching greedoids of Eulerian digraphs, the greedoid polynomial is related to the interior polynomial via duality. Let us sketch this connection. For rooted directed graphs, the greedoid polynomial of the induced branching greedoid is equivalent to the enumerator of graph parking functions [6] (also commonly called $G$-parking functions or generalized parking functions [19]) via a reversal of the sequence of the coefficients. Now, for Eulerian digraphs, the parking function enumerator agrees with the interior polynomial of the cographic matroid of the digraph [20]. Hence in this case the degree of the interior polynomial corresponds to the degree of the lowest term in the greedoid polynomial.
An edge set of a digraph is called a feedback arc set if it intersects each directed cycle. Using this dual notion of a dijoin, Theorem 1.6 implies the following.
Theorem 1.8.
The degree of the parking function enumerator of a connected Eulerian digraph $G$ (with any root) is equal to $|E(G)|-|V(G)|+1-{\rm minfas}(G)$, where ${\rm minfas}(G)$ denotes the minimal cardinality of a feedback arc set in $G$.
Equivalently, for the branching greedoid of $G$ (with any root), the coefficient of $x^{i}$ in the greedoid polynomial is zero for $i=0,\dots,{\rm minfas}(G)-1$, and nonzero for $i={\rm minfas}(G)$.
We remark that, in connected Eulerian digraphs, not just the degree but each coefficient of the parking function enumerator/greedoid polynomial is independent of the root [18, 6]. None of this holds for general digraphs.
We generalize Theorem 1.8 to branching greedoids of all rooted digraphs and, in a certain sense, to all greedoids. (These settings do not correspond to interior polynomials anymore.) To state our result on general directed branching greedoids, we define a rooted variant of a feedback arc set. For a graph $G=(V,E)$ and edge set $F\subseteq E$, we put $G[F]=(V,F)$. For a digraph $G$ and vertex $s$, we call the graph $s$-root-connected if each vertex is reachable along a directed path from $s$.
Definition 1.9.
Let $G$ be an $s$-root-connected digraph with edge set $E$. We say that a set of edges $F\subseteq E$ is an $s$-connected feedback arc set if $G[E-F]$ is an $s$-root-connected acyclic digraph. (Such a set $F$ always exists.) We denote by ${\rm minfas}(G,s)$ the minimum cardinality of an $s$-connected feedback arc set of $G$.
With that, our formula is as follows.
Theorem 1.10.
Let $G=(V,E)$ be an $r$-root-connected digraph. Then the degree of the parking function enumerator of $G$, rooted at $r$, is equal to $|E|-|V|+1-{\rm minfas}(G,r)$.
Equivalently, in the greedoid polynomial of the branching greedoid of $G$ rooted at $r$, the coefficients of $x^{0},\dots,x^{{\rm minfas}(G,r)-1}$ are zero, and the coefficient of $x^{{\rm minfas}(G,r)}$ is nonzero.
Remark 1.11.
By [5, Theorem 6.10], the constant term of the greedoid polynomial of a (root-connected) rooted digraph $G$ is zero if and only if $G$ contains a directed cycle. Theorem 1.10 strengthens this statement. Indeed, (for an $s$-root-connected digraph $G$) we have ${\rm minfas}(G,s)>0$ if and only if $G$ contains a directed cycle.
For general greedoids, we show that in order to determine the degree of the lowest term of the greedoid polynomial, it is in fact enough to understand when a greedoid has nonzero constant term in its greedoid polynomial.
Theorem 1.12.
Let $X=\{E,\mathcal{F}\}$ be a greedoid of rank $r$, and for a subset $S\subseteq E$ let $X|_{S}$ denote the greedoid obtained by restricting $X$ to $S$. Let
$k=\min\{|S|\mid S\subseteq E\text{ so that }{\rm rank}(X|_{E-S})=r\text{ and the constant term of the}\\ \text{greedoid polynomial of $X|_{E-S}$ is nonzero}\}.$
Then in the greedoid polynomial of $X$, the coefficient of $x^{i}$ is zero for $i=0,\dots,k-1$, and the coefficient of $x^{k}$ is nonzero.
In fact, Theorem 1.10 follows directly from Theorem 1.12 and the characterization of Björner, Korte, and Lovász mentioned in Remark 1.11.
The paper is structured as follows. Section 2 contains our results on the interior polynomial in the most ‘classical’ case of directed graphs, including the proofs of Theorems 1.1 and 1.3. Section 3 discusses the comparison of interior polynomials for different orientations of the same graph. In Section 4 we generalize Section 2 to regular oriented matroids and prove Theorems 1.6 and 1.7. Finally, in Section 5 we turn to parking function enumerators and greedoid polynomials, and prove Theorems 1.8, 1.12, and 1.10.
1.3. Acknowledgements
We are grateful to András Frank for stimulating conversations and for pointing out to us Theorem 9.6.12 of his book [8].
TK was supported by consecutive Japan Society for the Promotion of Science (JSPS) Grants-in-Aid for Scientific Research C (nos. 17K05244 and 23K03108).
LT was supported by the National Research, Development and Innovation Office of Hungary – NKFIH, grant no. 132488, by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences, and by the ÚNKP-22-5 New National Excellence Program of the Ministry for Innovation and Technology, Hungary. This work was also partially supported by the Counting in Sparse Graphs Lendület Research Group of the Alfréd Rényi Institute of Mathematics.
1.4. Graph notation
Let $G=(V,E)$ be a directed graph (abbreviated to digraph throughout). A non-empty set of edges $C^{*}\subseteq E(G)$ is called a cut if there is a partition $V_{0}\sqcup V_{1}=V(G)$ of the vertex set such that $C^{*}$ is the set of edges connecting a vertex of $V_{0}$ and a vertex of $V_{1}$. The sets $V_{0}$ and $V_{1}$ are called the shores of the cut. Note that in a connected graph, a cut uniquely determines its two shores. A cut $C^{*}$ is called elementary if it is minimal among cuts with respect to containment; this is equivalent to the condition that $V_{0}$ and $V_{1}$ both span (weakly) connected subgraphs. A cut $C^{*}$ is directed if either each edge of $C^{*}$ leads from $V_{0}$ to $V_{1}$, or each edge of $C^{*}$ leads from $V_{1}$ to $V_{0}$.
In a directed graph, we call a subgraph a spanning tree if it is a spanning tree (i.e., connected and cycle-free) when forgetting the orientations. For a spanning tree $T$ and an edge $e\notin T$, the fundamental cycle of $e$ with respect to $T$, denoted by $C(T,e)$, is the unique cycle in $T\cup e$. Similarly, for an edge $e\in T$, the fundamental cut of $e$ with respect to $T$ is the unique (elementary) cut in the complement of $T-e$. We denote it by $C^{*}(T,e)$.
A set of edges $K\subseteq E(G)$ is called a directed join or dijoin if for any directed cut $C^{*}$ of $G$ we have $C^{*}\cap K\neq\emptyset$. We denote the cardinality of a minimal dijoin of $G$ by $\nu(G)$. A set of edges $F\subseteq E(G)$ is called a feedback arc set if for any directed cycle $C$ we have $C\cap F\neq\emptyset$. We denote the minimal cardinality of a feedback arc set of $G$ by ${\rm minfas}(G)$.
Let $G$ be a digraph with a fixed root vertex $s$. The graph $G$ is said to be $s$-root-connected if each vertex is reachable along a directed path from $s$. We call a subgraph $A$ an arborescence rooted at $s$ if its underlying undirected graph is a tree and it is $s$-root-connected. A spanning arborescence is an arborescence that contains each vertex of $G$. We will denote the set of spanning arborescences of $G$, rooted at $s$, by ${\rm Arb}(G,s)$. We note that a digraph $G$, rooted at $s$, has a spanning arborescence (rooted at $s$) if and only if it is $s$-root-connected.
For a digraph $G=(V,E)$, and an edge set $S\subseteq E$, we denote by $G[S]$ the digraph with vertex set $V$ and edge set $S$.
2. The degree of the interior polynomial of a digraph
Let $G=(V,E)$ be a directed graph. To an edge $e=\overrightarrow{th}\in E$, let us associate the vector $\mathbf{x}_{e}=\mathbf{1}_{h}-\mathbf{1}_{t}\in\mathbb{R}^{V}$. (Here $t,h\in V$ and $\mathbf{1}_{t},\mathbf{1}_{h}\in\mathbb{R}^{V}$ are the corresponding generators.)
Definition 2.1.
The root polytope of a directed graph $G=(V,E)$ is the convex hull
$\mathcal{Q}_{G}=\operatorname{Conv}\{\,\mathbf{x}_{e}\mid e\in E\,\}\subset\mathbb{R}^{V}.$
The extended root polytope of $G$ is
$\tilde{\mathcal{Q}}_{G}=\operatorname{Conv}(\{\mathbf{0}\}\cup\{\,\mathbf{x}_{e}\mid e\in E\,\})\subset\mathbb{R}^{V}.$
Let us recall the well known facts that for $G=T$ a tree (or forest), both $\mathcal{Q}_{T}$ and $\tilde{\mathcal{Q}}_{T}$ are simplices, moreover, they are unimodular with respect to the lattice $\mathbb{Z}^{V}$. See, for example, [13] for details and proofs.
Since we are about to define the interior polynomial of a digraph as the $h^{*}$-polynomial of its extended root polytope, let us recall that for any $d$-dimensional polytope $Q\subset\mathbb{R}^{n}$ with vertices in $\mathbb{Z}^{n}$, its $h^{*}$-polynomial $\sum_{i=0}^{d}h^{*}_{i}t^{i}$, also commonly called the $h^{*}$-vector of $Q$, is defined by Ehrhart’s identity
(2.1) $\sum_{i=0}^{d}h^{*}_{i}t^{i}=(1-t)^{d+1}\mathrm{Ehr}_{Q}(t),\quad\text{where}\quad\mathrm{Ehr}_{Q}(t)=\sum_{k=0}^{\infty}|(k\cdot Q)\cap\mathbb{Z}^{n}|\,t^{k}$
is the so called Ehrhart series of $Q$. We note that $h^{*}_{0}=1$ whenever $d\geq 0$, i.e., whenever $Q$ is non-empty.
Intuitively, the $h^{*}$-polynomial can be thought of as a refinement of volume. Indeed, $h^{*}(1)$ (that is, the sum of the coefficients) is equal to the normalized volume of the polytope, where by normalized we mean that the volume of a $d$-dimensional unimodular simplex is $1$.
Now we are in a position to introduce our object of study for this section.
Definition 2.2 (Interior polynomial).
Let $G$ be a directed graph. We call the $h^{*}$-polynomial of the extended root polytope $\tilde{\mathcal{Q}}_{G}$ the interior polynomial of $G$, and denote it with $I_{G}$.
Remark 2.3.
In the earlier papers [10, 11, 13], the interior polynomial was defined only for so-called semi-balanced digraphs, and as the $h^{*}$-polynomial of $\mathcal{Q}_{G}$ instead of $\tilde{\mathcal{Q}}_{G}$. As we will soon point out, for these graphs, the $h^{*}$-polynomials of $\mathcal{Q}_{G}$ and $\tilde{\mathcal{Q}}_{G}$ agree. Thus our current definition contains the previous one.
Let us repeat our main claim about the interior polynomial.
Theorem 1.1.
Let $G=(V,E)$ be a connected digraph. The degree of the interior polynomial of $G$ is equal to $|V|-1-\nu(G)$, where
$\nu(G)=\min\{|K|\mid K\subseteq E\text{ is a dijoin of }G\}.$
Before proving Theorem 1.1, we examine root polytopes in more detail. If $G$ is connected, then the dimension of $\mathcal{Q}_{G}$ is either $|V|-1$ or $|V|-2$. It is $|V|-2$ if and only if $G$ satisfies the following condition [13].
Definition 2.4.
[13] A directed graph $G$ is semi-balanced if there is a function $\ell\colon V\to\mathbb{Z}$ such that we have $\ell(h)-\ell(t)=1$ for each edge $\overrightarrow{th}$ of $G$. We call such a function $\ell$ a layering of $G$.
We note that an alternative characterization of semi-balanced digraphs is that each cycle has the same number of edges going in the two directions around the cycle. This description was given as the definition in [13].
Now let us turn to the extended root polytope $\tilde{\mathcal{Q}}_{G}$. Note that if $G$ has a directed cycle $C$ then $\frac{1}{|C|}\sum_{e\in C}\mathbf{x}_{e}=\mathbf{0}$, that is, $\mathbf{0}$ is a point of $\mathcal{Q}_{G}$. Hence if $G$ is not acyclic, then $\mathcal{Q}_{G}=\tilde{\mathcal{Q}}_{G}$.
On the other hand, if $G$ is semi-balanced with layering $\ell$, then $\mathcal{Q}_{G}$ lies in the affine hyperplane $\{\mathbf{x}\in\mathbb{R}^{V}\mid\ell\cdot\mathbf{x}=1\}$. (Here we view the layering $\ell$ as a vector in $\mathbb{R}^{V}$ and use the standard dot product.) Thus, ${\tilde{\mathcal{Q}}}_{G}$ and $\mathcal{Q}_{G}$ do not coincide but as $\tilde{\mathcal{Q}}_{G}$ is just a cone (of the minimum possible height) over $\mathcal{Q}_{G}$, their $h^{*}$-vectors still agree.
Hence if we defined the interior polynomial using $\mathcal{Q}_{G}$ instead of the extended root polytope, that would only make a difference for acyclic, but not semi-balanced graphs. However, Theorem 1.1 would not hold if the $h^{*}$-polynomial of $\mathcal{Q}_{G}$ replaced that of ${\tilde{\mathcal{Q}}}_{G}$.
Example 2.5.
Consider the triangle $G=(\{v_{1},v_{2},v_{3}\},\{\overrightarrow{v_{1}v_{2}},\overrightarrow{v_{2}v_{3}},\overrightarrow{v_{1}v_{3}}\})$. It is easy to check that $h^{*}_{\mathcal{Q}_{G}}(x)=1$, while $h^{*}_{{\tilde{\mathcal{Q}}}_{G}}(x)=x+1$. Moreover we have $\nu(G)=1$, whence $|V(G)|-1-\nu(G)=1$. Thus indeed, $h^{*}_{\mathcal{Q}_{G}}$ does not satisfy the degree formula, only $h^{*}_{{\tilde{\mathcal{Q}}}_{G}}$ does.
As an alternative way to think of the extended root polytope, note that loop edges $f$ have $\mathbf{x}_{f}=\mathbf{0}$. Hence ${\tilde{\mathcal{Q}}}_{G}=\mathcal{Q}_{G\cup f}$, where $f$ is a loop edge attached to any vertex of $G$.
We will need a description of $\tilde{\mathcal{Q}}_{G}$ by linear inequalities (half-spaces). That requires the following notions.
Definition 2.6.
Let $C^{*}$ be a cut in the graph $G$ with shores $V_{0}$ and $V_{1}$. Let $f_{C^{*}}$ be the linear functional with $f_{C^{*}}(\mathbf{1}_{v})=1$ when $v\in V_{1}$ and $f_{C^{*}}(\mathbf{1}_{v})=0$ when $v\in V_{0}$. If $G$ is directed and $C^{*}$ is a directed cut, we will always suppose that $V_{1}$ is the shore containing the heads of the edges in the cut. We will refer to $f_{C^{*}}$ as the functional induced by the cut $C^{*}$.
Definition 2.7.
Let $G=(V,E)$ be a directed graph. A function $\ell\colon V\to\mathbb{Z}$ is called an admissible layering of $G$ if $\ell(h)-\ell(t)\leq 1$ holds for each $\overrightarrow{th}\in E$, and the edges with $\ell(h)-\ell(t)=1$ form a (weakly) connected subgraph of $G$.
We remark that any admissible layering of a semi-balanced graph is actually a layering in the sense of Definition 2.4, so we have not introduced anything new into that context.
Then, our description is as follows.
Proposition 2.8.
The extended root polytope of any connected digraph $G$ satisfies
$\tilde{\mathcal{Q}}_{G}=\left\{\mathbf{x}\in\mathbb{R}^{V}\middle|\begin{array}[]{cl}f_{C^{*}}(\mathbf{x})\geq 0&\text{for all elementary directed cuts }C^{*}\text{ of }G\\ \ell\cdot\mathbf{x}\leq 1&\text{for all admissible layerings $\ell$ of $G$}\\ \mathbf{1}\cdot\mathbf{x}=0&\end{array}\right\}.$
Here $\mathbf{1}=\sum_{v\in V}\mathbf{1}_{v}\in\mathbb{R}^{V}$. Note that if $\ell$ is an admissible layering then so is $\ell+m\cdot\mathbf{1}$ for all $m\in\mathbb{Z}$. As is obvious from the Proposition, members of such equivalence classes of admissible layerings describe the same constraint for $\tilde{\mathcal{Q}}_{G}$. If we chose a representative of each class, for example by assuming that $\ell(v_{0})=0$ for some fixed vertex $v_{0}$, then the number of constraints would become formally finite. (Indeed, there are only finitely many ways to choose the connected subgraph whose edges $\overrightarrow{th}$ satisfy $\ell(h)-\ell(t)=1$, and with $\ell(v_{0})=0$ that subgraph already determines $\ell$.)
Proposition 2.8 contains some previously published special cases, such as [11, Proposition 3.6] and [9, Theorem 3.1]. A very similar facet description of $\mathcal{Q}_{G}$ was recently obtained by Numata et al. [17].
Proof of Proposition 2.8.
First note that each vector $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$ satisfies the conditions of the right hand side. Clearly $\mathbf{1}\cdot\mathbf{x}_{e}=0$ for each $e\in E(G)$ and $\mathbf{1}\cdot\mathbf{0}=0$, whence $\mathbf{1}\cdot\mathbf{x}=0$ for each $\mathbf{x}\in\tilde{\mathcal{Q}}_{G}$. Similarly for any admissible layering $\ell$, by definition we have $\ell\cdot\mathbf{x}_{e}\leq 1$ for each $e\in E$, which in addition to the obvious $\ell\cdot\mathbf{0}=0$ implies $\ell\cdot\mathbf{x}\leq 1$ for each $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$. Finally let $C^{*}$ be any directed cut. Then by definition, the induced functional $f_{C^{*}}$ is such that $f_{C^{*}}(\mathbf{x}_{e})=0$ for $e\notin C^{*}$ and $f_{C^{*}}(\mathbf{x}_{e})=1$ for $e\in C^{*}$; moreover $f_{C^{*}}(\mathbf{0})=0$. Hence each $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$ satisfies $f_{C^{*}}(\mathbf{x})\geq 0$.
Conversely, we show that any element $\mathbf{x}$ of the right hand side belongs to ${\tilde{\mathcal{Q}}}_{G}$. From the first part of the proof, and the fact that ${\tilde{\mathcal{Q}}}_{T}$ is a $(|V|-1)$-dimensional simplex for any spanning tree $T\subseteq G$, it is already clear that $\dim{\tilde{\mathcal{Q}}}_{G}=|V|-1$. The condition $\mathbf{1}\cdot\mathbf{x}=0$ says that $\mathbf{x}$ is in the same hyperplane of $\mathbb{R}^{V}$ as ${\tilde{\mathcal{Q}}}_{G}$; in the rest of the proof we work relative to this hyperplane.
Consider an arbitrary facet $F$ of ${\tilde{\mathcal{Q}}}_{G}$. It suffices to show that $\mathbf{x}$ lies either in the hyperplane of $F$ or on the same side of the hyperplane as the interior of ${\tilde{\mathcal{Q}}}_{G}$.
The facet $F$ needs to contain $|V|-1$ affine independent vertices of ${\tilde{\mathcal{Q}}}_{G}$. If $F$ does not contain $\mathbf{0}$, then those are $|V|-1$ affine independent vertices of type $\mathbf{x}_{e}$, which means that the corresponding edges $e$ form a spanning tree $T$. The tree $T$, along with the orientations of its edges, determines a layering $\ell\colon V\to\mathbb{Z}$ so that $\ell(\mathbf{x}_{e})=1$ for all $e\in T$. This layering is necessarily admissible, for otherwise there would be an edge $f$ with $\ell(\mathbf{x}_{f})\geq 2$, meaning that the vertices $\mathbf{0}$ and $\mathbf{x}_{f}$ lie on opposite sides of the hyperplane of $F$. Therefore the hyperplane of the facet is $\{\mathbf{x}\mid\ell\cdot\mathbf{x}=1\}$ for the admissible layering $\ell$, and as $\ell\cdot\mathbf{0}=0$, the polytope $\tilde{\mathcal{Q}}_{G}$ lies in the half-space $\{\mathbf{p}\mid\ell\cdot\mathbf{p}\leq 1\}$, whence indeed the hyperplane of $F$ does not separate $\mathbf{x}$ from ${\tilde{\mathcal{Q}}}_{G}$.
If $F$ contains $\mathbf{0}$, then it must additionally contain $|V|-2$ affine independent vectors of the form $\mathbf{x}_{e}$. Here the corresponding edges $e$ form a two-component forest in $G$, whose components give rise to an elementary cut $C^{*}$. That in turn induces the functional $f_{C^{*}}$ that vanishes along $F$. This and $F$ being a facet imply that $C^{*}$ is a directed cut, for otherwise there would exist vectors $\mathbf{x}_{e}$ on both sides of the kernel of $f_{C^{*}}$. As the vectors $\mathbf{x}_{e}$ for $e\in C^{*}$ have $f_{C^{*}}(\mathbf{x}_{e})=1$, clearly ${\tilde{\mathcal{Q}}}_{G}$ lies in the half-space $\{\mathbf{p}\mid f_{C^{*}}(\mathbf{p})\geq 0\}$. Thus again, the hyperplane of $F$ does not separate $\mathbf{x}$ from ${\tilde{\mathcal{Q}}}_{G}$. ∎
Note that we can also conclude from the previous proof that (within the subspace $\{\mathbf{x}\mid\mathbf{1}\cdot\mathbf{x}=0\}$), the hyperplanes $\{\mathbf{x}\mid f_{C^{*}}(\mathbf{x})=0\}$ for elementary directed cuts $C^{*}$ and the hyperplanes $\{\mathbf{x}\mid\ell\cdot\mathbf{x}=1\}$ for admissible layerings $\ell$ all contain facets of ${\tilde{\mathcal{Q}}}_{G}$, and collectively these are all the facets of ${\tilde{\mathcal{Q}}}_{G}$.
A key ingredient in proving Theorem 1.1 will be the following corollary of Ehrhart–Macdonald reciprocity.
Theorem 2.9.
[3, Theorem 4.5] Let $P\subset\mathbb{R}^{n}$ be a $d$-dimensional ($d\geq 0$) lattice polytope with $h^{*}$-polynomial $h^{*}_{d}t^{d}+\dots+h^{*}_{1}t+1$. Then $h^{*}_{d}=\dots=h^{*}_{k+1}=0$ and $h^{*}_{k}\neq 0$ if and only if $(d-k+1)P$ is the smallest integer dilate of $P$ that contains a lattice point in its relative interior.
Notice here that $(d+1)P$ certainly contains an interior lattice point. The degree of the $h^{*}$-polynomial of $P$ tells us exactly ‘how much sooner’ such a point occurs.
In our cases, $\tilde{\mathcal{Q}}_{G}$ is a $(|V|-1)$-dimensional polytope. Thus if we show that
$(\nu(G)+1)\cdot\tilde{\mathcal{Q}}_{G}$
is the smallest integer dilate of the extended root polytope that contains a lattice point in its interior, then by Theorem 2.9, it follows that the degree of $I_{G}$ is indeed $|V|-1-\nu(G)$. For this, we need to clarify the connection between dijoins of $G$ and interior points of ${\tilde{\mathcal{Q}}}_{G}$.
First, let us make a simple observation about dijoins. Let $C$ be any cycle in $G$. We call $C$ an signed cycle if we have picked one of its cyclic orientations as “positive.” Then we can partition $C$ to $C^{-}\sqcup C^{+}$ where $C^{-}$ is the set of edges pointing in the negative cyclic direction, and $C^{+}$ is the set of edges pointing in the positive cyclic direction. Signed cycles are not to be confused with directed cycles; the latter means a cycle whose edges all point in the same cyclic direction.
Suppose that $C$ is a signed cycle with a choice of orientation such that $|C^{-}|>|C^{+}|$. If a dijoin $K$ contains $C^{-}$, then $K$ is not minimal. Indeed, in this case $|K-C^{-}\cup C^{+}|<|K|$ and $K-C^{-}\cup C^{+}$ is also a dijoin. This is so because if a directed cut does not intersect $C$, then it is covered by $K-C^{-}\cup C^{+}$ the same way as by $K$; if a directed cut intersects $C$ then it intersects both $C^{-}$ and $C^{+}$, hence it is also covered by $K-C^{-}\cup C^{+}$.
Let us call a dijoin $K$ a long arc dijoin if it contains $C^{-}$ for some signed cycle $C$ with $|C^{-}|>|C^{+}|$. Our previous observation implies that long arc dijoins are never minimal. Note also that $C^{-}$ and $C^{+}$ are not meant to be actual arcs of $C$.
We claim the following characterization of relative interior points of ${\tilde{\mathcal{Q}}}_{G}$. The distinction between $<1$ and $\leq 1$ is a crucial subtlety in the statement. Note that since $\mathbf{0}$ is a generator of ${\tilde{\mathcal{Q}}}_{G}$, our formulas can be thought of as convex linear combinations even when the sum of the coefficients is less than $1$.
Proposition 2.10.
A point $\mathbf{p}\in\tilde{\mathcal{Q}}_{G}$ is in the relative interior of ${\tilde{\mathcal{Q}}}_{G}$ if and only if there exists a cycle-free dijoin $K$ of $G$ such that $\mathbf{p}=\sum_{e\in K}\lambda_{e}\mathbf{x}_{e}$, where $\lambda_{e}>0$ for each $e\in K$ and $\sum_{e\in K}\lambda_{e}<1$.
Proof.
By Proposition 2.8, a point $\mathbf{p}\in{\tilde{\mathcal{Q}}}_{G}$ is in the interior of $\tilde{\mathcal{Q}}_{G}$ if and only if $f_{C^{*}}(\mathbf{p})>0$ for each directed cut $C^{*}$ and $\ell\cdot\mathbf{p}<1$ for each admissible layering $\ell$. Recall that the functional induced by $C^{*}$ satisfies $f_{C^{*}}(\mathbf{x}_{e})=0$ whenever $e\notin C^{*}$, and $f_{C^{*}}(\mathbf{x}_{e})=1$ for each $e\in C^{*}$.
We start with the ‘if’ direction. Suppose that $K$ is a dijoin and $\mathbf{p}=\sum_{e\in K}\lambda_{e}\mathbf{x}_{e}$, with $\sum_{e\in K}\lambda_{e}<1$ and $\lambda_{e}>0$ for each $e\in K$. (The cycle-freeness of $K$ is not needed for this direction.) Then $\mathbf{p}=\sum_{e\in K}\lambda_{e}\mathbf{x}_{e}+(1-\sum_{e\in K}\lambda_{e})\cdot\mathbf{0}$, in particular $\mathbf{p}\in\tilde{\mathcal{Q}}_{G}$. For any directed cut $C^{*}$, we have $f_{C^{*}}(\mathbf{p})=\sum_{e\in C^{*}\cap K}\lambda_{e}>0$, since the intersection is nonempty (by the definition of a dijoin), and the summands are all positive. Similarly, for any admissible layering $\ell$ we have $\ell\cdot\mathbf{x}_{e}\leq 1$ for any $e\in E$, and $\ell\cdot\mathbf{0}=0$. Since the coefficient of $\mathbf{0}$ is positive, this implies $\ell\cdot\mathbf{p}<1$. Thus $\mathbf{p}$ is in the interior of ${\tilde{\mathcal{Q}}}_{G}$.
Now we turn to the ‘only if’ direction. Let $\mathbf{p}$ be an arbitrary point in the relative interior of ${\tilde{\mathcal{Q}}}_{G}$, and take any convex combination $\mathbf{p}=\sum_{e\in S}\lambda_{e}\mathbf{x}_{e}+\mu\cdot\mathbf{0}$ with $\lambda_{e}>0$ for all $e\in S$, where $S\subseteq E$ is some subset. For a point of ${\tilde{\mathcal{Q}}}_{G}$, we can always find such a formula (usually more than one). We claim that if $\mathbf{p}$ is in the interior, then $S$ is necessarily a dijoin. Indeed, suppose that $S$ is disjoint from a directed cut $C^{*}$. Then $f_{C^{*}}(\mathbf{p})=\sum_{e\in S}\lambda_{e}\cdot f_{C^{*}}(\mathbf{x}_{e})+\mu\cdot f_{C^{*}}(\mathbf{0})=0$, which contradicts $\mathbf{p}$ being an interior point of ${\tilde{\mathcal{Q}}}_{G}$.
Next we show that either $\sum_{e\in S}\lambda_{e}<1$ or $S$ is a long-arc dijoin. After that, we will establish that if $S$ is a long-arc dijoin, then there exists another dijoin $S^{\prime}$ such that $\mathbf{p}=\sum_{e\in S^{\prime}}\lambda^{\prime}_{e}\mathbf{x}_{e}$ and $\sum_{e\in S^{\prime}}\lambda^{\prime}_{e}<1$. Finally, we will show how to remove edges from the dijoin to ascertain cycle-freeness.
So suppose that the dijoin $S$ we found for $\mathbf{p}$ is such that $\sum_{e\in S}\lambda_{e}=1$. Assume for a contradiction that $S$ is not a long-arc dijoin, that is, for all signed cycles $C$ so that $|C^{-}|>|C^{+}|$, it does not contain $C^{-}$. We will reach a contradiction by showing that $\mathbf{p}$ is on the boundary of ${\tilde{\mathcal{Q}}}_{G}$. More concretely, we will find a vector $\ell$ such that $\ell\cdot\mathbf{x}_{e}=1$ for each $e\in S$ and $\ell\cdot\mathbf{x}_{e}\leq 1$ for each $e\notin S$. The existence of such a vector follows easily from the Farkas lemma and we will rely on that argument in Section 4 in the regular matroid case. Here however, let us construct $\ell$ in a more elementary way.
Choose a spanning forest within $G[S]$ and use it to fix integer values $\ell(v)$ for each $v\in V$ in such a way that $\ell\cdot\mathbf{x}_{e}=1$ for each edge $e$ in the forest. This can be done uniquely, up to one additive constant for each connected component of $G[S]$. For other edges of $S$ within one component, $\ell\cdot\mathbf{x}_{e}=1$ is automatically satisfied because they make fundamental cycles $C\subseteq S$ with the forest and such cycles have to have $|C^{-}|=|C^{+}|$ by our assumption that $S$ is not a long-arc dijoin.
It remains to ensure that $\ell\cdot\mathbf{x}_{e}\leq 1$ for edges $e\in E-S$. For some choice of $\ell$ as above, let us consider the set of ‘bad’ edges that violate this condition, that is
$B_{\ell}=\{e\in E-S\mid\ell\cdot\mathbf{x}_{e}>1\},\text{ as well as the set }A_{\ell}=\{e\in E-S\mid\ell\cdot\mathbf{x}_{e}=1\}$
where a violation ‘almost’ occurs. We use $B_{\ell}\cup A_{\ell}$ as the set of edges in an auxiliary graph $G^{\prime}$, in which the vertices are the connected components of $G[S]$. (I.e., $G^{\prime}$ is the graph obtained by contracting each connected component of $G[S]$ to a point and only keeping elements of $B_{\ell}$ and $A_{\ell}$ as edges, some of which may become loops.)
Suppose that $C^{\prime}$ is a directed cycle in $G^{\prime}$. Then there exists a pullback cycle $C$ in $G$ that is obtained by adding edges of $S$ to $C^{\prime}$. As $C^{\prime}$ is directed, $C$ can be oriented in such a way that $C^{\prime}\subseteq C^{+}$, and hence $C^{-}\subseteq S$. Because $\ell$ jumps by at least $1$ along each edge of $C^{\prime}$ and by exactly $1$ along edges of $S$, we have to have $|C^{-}|\geq|C^{+}|$. But since $C^{-}\subseteq S$ and $S$ is not a long-arc dijoin, only $|C^{-}|=|C^{+}|$ is possible and that implies $C^{\prime}\subseteq A_{\ell}$.
Now as elements of $B_{\ell}$ cannot be in a directed cycle in $G^{\prime}$, they all must be in directed cuts of $G^{\prime}$. For any such cut, we may lower the values of $\ell$ by $1$ over all connected components of $G[S]$ that form the head-shore of the cut. It is clear from the construction that this new vector $\ell^{\prime}$ has $B_{\ell^{\prime}}\subseteq B_{\ell}$, moreover
$\sum_{e\in B_{\ell^{\prime}}}\ell^{\prime}\cdot\mathbf{x}_{e}<\sum_{e\in B_{\ell}}\ell\cdot\mathbf{x}_{e}.$
We may repeat our modification step for $\ell^{\prime}$ instead of $\ell$, and so on as long as $B$ is non-empty, but the last observation ensures that the process will be finite. At the end of it $B$ does become empty and we have found the desired vector $\ell$ such that $\ell\cdot\mathbf{x}_{e}\leq 1$ for $e\in E-S$ and $\ell\cdot\mathbf{x}_{e}=1$ for $e\in S$. This contradicts the assumption that $\mathbf{p}$ is an interior point. Hence we have proved that $\sum_{e\in S}\lambda_{e}=1$ implies that $S$ is a long-arc dijoin.
Next, we show that if $\mathbf{p}=\sum_{e\in S}\lambda_{e}\mathbf{x}_{e}$ with $\lambda_{e}>0$ for all $e\in S$ and $\sum_{e\in S}\lambda_{e}=1$, where $S$ is a long-arc dijoin, then we can find a different dijoin $S^{\prime}$ such that $\mathbf{p}=\sum_{e\in S^{\prime}}\lambda^{\prime}_{e}\mathbf{x}_{e}$ with $\lambda^{\prime}_{e}>0$ for all $e\in S^{\prime}$ and $\sum_{e\in K}\lambda^{\prime}_{e}<1$.
Let $C$ be a signed cycle that has $C^{-}\subseteq S$ and $|C^{-}|>|C^{+}|$. Such a cycle exists because of the definition of a long-arc dijoin. Let $\delta=\min\{\lambda_{e}\mid e\in C^{-}\}$ and let $M=\{e\in C^{-}\mid\lambda_{e}=\delta\}$. Then, by the obvious $\sum_{e\in C^{-}}\mathbf{x}_{e}=\sum_{e\in C^{+}}\mathbf{x}_{e}$, we have $\mathbf{p}=\sum_{e\in S-M\cup C^{+}}\lambda^{\prime}_{e}\mathbf{x}_{e}+\mu^{\prime}\cdot\mathbf{0}$ where
$\lambda^{\prime}_{e}=\left\{\begin{array}[]{cl}\lambda_{e}&\text{if $e\notin C$},\\ \lambda_{e}-\delta&\text{if $e\in C^{-}$},\\ \lambda_{e}+\delta&\text{if $e\in C^{+}$},\end{array}\right.$
and $\mu^{\prime}=(|C^{-}|-|C^{+}|)\cdot\delta$. This is a new convex combination for $\mathbf{p}$, with positive coefficients. Moreover, $S-M\cup C^{+}$ is also a dijoin because any directed cut that intersects $C$ necessarily intersects both $C^{+}$ and $C^{-}$. Hence we have found a linear combination for $\mathbf{p}$ using a dijoin and $\mathbf{0}$, such that the coefficient of $\mathbf{0}$ is positive.
It is left to show that $S$ can be chosen cycle-free. If there is any cycle $C\subseteq S$, then choose its orientation so that $|C^{-}|\geq|C^{+}|$ and apply the same procedure as in the previous paragraph. That is, lower the coefficients on elements of $C^{-}$ and raise them on elements of $C^{+}$, all by the same amount, until some coefficients become 00. Remove the corresponding edges from $S$. This does not increase the sum of the coefficients so it remains strictly less than $1$. The remaining edges still form a dijoin. The value of the linear combination does not change, i.e., it remains $\mathbf{p}$. We may continue this until the dijoin becomes cycle-free. ∎
The following Lemma is equivalent to the well-known fact that ${\tilde{\mathcal{Q}}}_{F}$ is a unimodular simplex for any forest $F$. See, e.g., [13, Corollary 3.6] for a proof.
Lemma 2.11.
For a forest $F$ and any positive integer $s$, a point $\mathbf{p}\in s\cdot{\tilde{\mathcal{Q}}}_{F}$ is a lattice point if and only if $\mathbf{p}=\sum_{e\in F}\mu_{e}\mathbf{x}_{e}$, where each $\mu_{e}$ are integer.
Proof of Theorem 1.1.
Let $K$ be a dijoin of $G$ with cardinality $\nu(G)$. Then $\mathbf{p}=\sum_{e\in K}\mathbf{x}_{e}+\mathbf{0}$ is a point of $(\nu(G)+1)\cdot\tilde{\mathcal{Q}}_{G}$, moreover, it clearly has integer coordinates. Now by Proposition 2.10 we have that $\mathbf{q}=\frac{1}{\nu(G)+1}\mathbf{p}=\sum_{e\in K}\frac{1}{\nu(G)+1}\mathbf{x}_{e}$ is an interior point of $\tilde{\mathcal{Q}}_{G}$, which implies that $\mathbf{p}$ is also an interior point of $(\nu(G)+1)\cdot\tilde{\mathcal{Q}}_{G}$.
We also need to prove that for $s\leq\nu(G)$, there is no interior lattice point in $s\cdot\tilde{\mathcal{Q}}_{G}$. Suppose that there is an interior lattice point $\mathbf{p}\in s\cdot\tilde{\mathcal{Q}}_{G}$ for some $s\in\mathbb{Z}_{>0}$ and consider $\mathbf{q}=\frac{1}{s}\mathbf{p}$, which is an interior point of $\tilde{\mathcal{Q}}_{G}$. Then by Proposition 2.10 there is a cycle-free dijoin $K$ such that $\mathbf{q}=\sum_{e\in K}\lambda_{e}\mathbf{x}_{e}$, where $\lambda_{e}>0$ for each $e\in K$, and $\sum_{e\in K}\lambda_{e}<1$.
Now we may apply Lemma 2.11 to $s$, $K$, and $\mathbf{p}=\sum_{e\in K}s\lambda_{e}\mathbf{x}_{e}$. This tells us that for $\mathbf{p}$ to be an integer vector, $s\lambda_{e}$ needs to be a positive integer for each $e\in K$. Hence altogether, we have $s>\sum_{e\in K}s\lambda_{e}\geq|K|\geq\nu(G)$. ∎
Next, to compute the leading coefficient, that is to prove Theorem 1.3, we use another result (a natural continuation of Theorem 2.9) from Ehrhart theory.
Theorem 2.12.
[2, Remark 1.2] Let $P$ be a $d$-dimensional lattice polytope with $h^{*}$-polynomial $h^{*}(t)=h^{*}_{d}t^{d}+\dots+h^{*}_{1}t+h^{*}_{0}$ such that $h^{*}_{d}=\dots=h^{*}_{k+1}=0$ but $h^{*}_{k}\neq 0$. Then $h^{*}_{k}$ equals to the number of lattice points in the relative interior of $(d-k+1)P$.
Hence to compute the leading coefficient of $I_{G}$, we need to characterize the interior lattice points in $(\nu(G)+1)\cdot{\tilde{\mathcal{Q}}}_{G}$. Such a result can be easily deduced from Proposition 2.10 and Lemma 2.11.
Proposition 2.13.
The interior lattice points of $(\nu(G)+1)\cdot{\tilde{\mathcal{Q}}}_{G}$ are exactly the net degree vectors of the minimal dijoins of $G$.
Proof.
By Proposition 2.10, $\mathbf{q}$ is an interior point of $(\nu(G)+1)\cdot{\tilde{\mathcal{Q}}}_{G}$ if and only if there is a cycle-free dijoin $K$ of $G$ such that $\mathbf{q}=\sum_{e\in K}\lambda_{e}\mathbf{x}_{e}$ where $\lambda_{e}>0$ for each $e\in K$ and $\sum_{e\in E}\lambda_{e}<\nu(G)+1$.
By Lemma 2.11, such a $\mathbf{q}$ is a lattice point if and only if each $\lambda_{e}$ is integer. Hence we conclude that $\lambda_{e}=1$ for each $e\in K$ and $|K|\leq\nu(G)$. Consequently $\mathbf{q}$ is an interior lattice point if and only if $K$ is a minimum cardinality dijoin, and $\mathbf{q}$ is the net degree vector of $K$. ∎
Proof of Theorem 1.3.
The statement follows at once from Theorem 1.1, Proposition 2.13, and Theorem 2.12. ∎
3. Degree-minimizing orientations
Let $\mathcal{G}$ be an undirected graph. It is natural to ask about the relationship between interior polynomials of different orientations of $\mathcal{G}$.
In [12] we looked at a special case of this problem, and considered the semi-balanced orientations of a bipartite graph $\mathcal{G}$ with partite classes $U$ and $W$. Any bipartite graph $\mathcal{G}$ has some (typically, many) semi-balanced orientations, but there are two special ones among them: The one where each edge is oriented from $U$ to $W$, and the one where each edge is oriented from $W$ to $U$. It is easy to see that these orientations are indeed semi-balanced. We call them the standard orientations of $\mathcal{G}$. The root polytopes of the two standard orientations are isometric, as they are reflections of each other. In particular, their interior polynomials coincide.
After looking at several examples, in [12] we conjectured that among all semi-balanced orientations of $\mathcal{G}$, the standard orientations minimize every coefficient of the interior polynomial. See [13, Example 6.5] for some concrete instances of this phenomenon. With Theorem 1.1 in hand, we are able to prove a weakened version of the conjecture, showing that the degree of the interior polynomial is minimized by the standard orientations. In fact, we can prove this among all orientations.
Theorem 3.1.
Let $\mathcal{G}$ be a connected, undirected bipartite graph with partite classes $U$ and $W$. Then among all orientations of $\mathcal{G}$, the degree of the interior polynomial is minimized by the standard orientations.
Proof.
Let $G$ be an arbitrary semi-balanced orientation of $\mathcal{G}$. By the Lucchesi–Younger theorem, the maximal number of disjoint directed cuts in $G$ is equal to $\nu(G)$, the minimal cardinality of a dijoin in $G$.
As the degree of $I_{G}$ is equal to $|V(G)|-1-\nu(G)$, the degree of $I_{G}$ is minimized for those orientations of $\mathcal{G}$ that maximize the number of disjoint directed cuts.
Let $c(\mathcal{G})$ denote the maximal number of disjoint cuts in $\mathcal{G}$. Clearly, for any orientation $G$ of $\mathcal{G}$, we have $\nu(G)\leq c(\mathcal{G})$, since if we take $\nu(G)$ disjoint directed cuts in $G$, those correspond to disjoint cuts in $\mathcal{G}$.
On the other hand, [8, Theorem 9.6.12] claims that the standard orientations have $c(\mathcal{G})$ disjoint directed cuts. Hence they do maximize $\nu$ among all orientations of $\mathcal{G}$, as claimed. ∎
Concrete examples (e.g., [13, Example 6.5]) show that typically, there are some non-standard orientations that also minimize the degree of the interior polynomial. It would be interesting to give a characterization for all the other orientations that attain the minimal degree.
For non-bipartite graphs, we have no candidates for the degree-minimizing orientations. Yet, based on computer calculations, we make the following conjecture.
Conjecture 3.2.
For any undirected graph, there exists an orientation that coefficientwise minimizes the $h^{*}$-polynomial of the extended root polytope among all orientations. Also, there exists an orientation that coefficientwise maximizes the $h^{*}$-polynomial of the extended root polytope among all orientations.
4. A generalization to regular matroids
It turns out that many properties of the root polytopes of digraphs extend word-by-word to regular oriented matroids, moreover, in some applications, one needs this more general case (see, for example, Section 5 or [20]). Here we briefly recall the notion of a regular matroid, along with ways of orienting such a structure, and then generalize the results of Section 2 to this context.
From among the many equivalent characterizations of the class of regular matroids, we will use the following: A matroid is regular if it can be represented by the column vectors of a totally unimodular matrix. Here a matrix is totally unimodular if each of its subdeterminants is either 00, $-1$, or $1$.
More concretely, let $A$ be an $n\times m$ totally unimodular matrix with columns $\{\mathbf{a}_{i}\}_{i\in E}$. The ground set of our matroid is going to be $E$. A set of elements $C=\{i_{1},\dots,i_{s}\}\subseteq E$ is a circuit if the corresponding columns $\mathbf{a}_{i_{1}},\dots,\mathbf{a}_{i_{s}}$ are minimally linearly dependent (over $\mathbb{R}$). Total unimodularity implies that in this case, the coefficients of the linear relation can be chosen from $\{-1,1\}$. For a circuit $C$, we call the two vectors in $\ker(A)\cap\{0,1,-1\}^{n}$, whose support is $C$, the vectors of $C$.
The orientation of the matroid is defined as follows. For each circuit $C$, any vector $\chi_{C}$ of $C$ gives a partition $C=C^{+}\sqcup C^{-}$, where $C^{+}$ is the subset of $C$ with positive coordinates in $\chi_{C}$, and $C^{-}$ is the subset of $C$ with negative coordinates. For different vectors of $C$, the subsets $C^{+}$ and $C^{-}$ might switch roles but the partition itself is well-defined because of the minimality of $C$. By a regular oriented matroid we mean a pair $(E,\mathcal{C})$, where $\mathcal{C}$ is the system of (unordered) pairs $(C^{+},C^{-})$ defined as above from a totally unimodular matrix $A$.
By a signed circuit, we mean a circuit $C$ with a choice of orientation, that is, with a choice of a vector of $C$ that determines which subset of $C$ should be $C^{+}$ and which one should be $C^{-}$.
Let us stress that a given regular oriented matroid may have several different totally unimodular representing matrices. (For example, row elimination steps performed on $A$ do not change the oriented matroid.)
One important example of regular oriented matroids is that of directed graphs. To a directed graph, one can associate a regular oriented matroid using its (directed) vertex-edge incidence matrix. The circuits of this matroid correspond to the cycles of the graph, and the partition $C=C^{+}\sqcup C^{-}$ of a circuit is to edges of the cycle pointing in the two cyclic directions.
We call a subset $C^{*}\subseteq E$ a cocircuit if there is a linear functional $h\colon\mathbb{R}^{n}\to\mathbb{R}$, with kernel the hyperplane $H\subseteq\mathbb{R}^{n}$, such that $\mathbf{a}_{k}\in H$ if $k\notin C^{*}$ and $\mathbf{a}_{k}\notin H$ if $k\in C^{*}$; moreover, $C^{*}$ is minimal with respect to this property. Cocircuits generalize elementary cuts of graphs. By the total unimodularity of $A$, we may suppose that $h(\mathbf{a}_{k})\in\{0,1,-1\}$ for each $k\in E$. We say that $k\in(C^{*})^{+}$ if $h(\mathbf{a}_{k})=1$ and $k\in(C^{*})^{-}$ if $h(\mathbf{a}_{k})=-1$. Again, $(C^{*})^{+}$ and $(C^{*})^{-}$ can switch roles, but the partition $C^{*}=(C^{*})^{+}\sqcup(C^{*})^{-}$ is well-defined.
A cocircuit is called directed if either $(C^{*})^{+}$ or $(C^{*})^{-}$ is empty. In this case we will always suppose that $(C^{*})^{-}$ is the empty part. A set $K\subseteq E$ is called a dijoin if $K$ intersects each directed cocircuit.
We call a regular oriented matroid co-Eulerian if $|C^{+}|=|C^{-}|$ for each circuit $C$. For graphic oriented matroids, being co-Eulerian is equivalent to the orientation of the graph being semi-balanced.
If $A$ is a totally unimodular matrix with columns $\mathbf{a}_{1},\dots,\mathbf{a}_{m}$, then its root polytope is defined as $\mathcal{Q}_{A}=\operatorname{Conv}\{\mathbf{a}_{1},\dots,\mathbf{a}_{m}\}$, and the extended root polytope is
$\tilde{\mathcal{Q}}_{A}=\operatorname{Conv}\{\mathbf{0},\mathbf{a}_{1},\dots,\mathbf{a}_{m}\}.$
It turns out that if $A$ and $A^{\prime}$ are two totally unimodular matrices representing the same oriented matroid $M$, then the $h^{*}$-polynomials of ${\tilde{\mathcal{Q}}}_{A}$ and ${\tilde{\mathcal{Q}}}_{A^{\prime}}$ are the same [20]. (We note that [20] proves this for $\mathcal{Q}_{A}$ and $\mathcal{Q}_{A^{\prime}}$, but a straightforward modification of the argument yields the result also for ${\tilde{\mathcal{Q}}}_{A}$ and ${\tilde{\mathcal{Q}}}_{A^{\prime}}$.) Hence the $h^{*}$-polynomial of ${\tilde{\mathcal{Q}}}_{A}$ is an invariant of the regular oriented matroid $M$, which we call the interior polynomial and denote by $I_{M}$. Note that the orientation of the matroid matters: if we keep the (unoriented) matroid structure, but change the orientation, then the interior polynomial might change. (This is true even for graphs, cf. [13, Example 6.5].)
Theorem 1.6.
Let $M$ be a regular oriented matroid of rank $r$. Then the degree of $I_{M}$ is equal to $r-\nu(M)$, where
$\nu(M)=\min\{|K|\mid K\subseteq E\text{ is a dijoin of }M\}.$
The proof proceeds through the same steps as in the graph case. First, we again need a facet description for ${\tilde{\mathcal{Q}}}_{A}$.
Let us start by examining how facets of ${\tilde{\mathcal{Q}}}_{A}$ not containing the origin look. Take a maximal affine independent set of vectors $\mathbf{a}_{i_{1}},\dots,\mathbf{a}_{i_{s}}$ along such a facet. Then together with $\mathbf{0}$ they form a maximal affine independent set among the generators of ${\tilde{\mathcal{Q}}}_{A}$. This happens if and only if $\mathbf{a}_{i_{1}},\dots,\mathbf{a}_{i_{s}}$ are a maximal linearly independent set, that is, the corresponding elements form a basis in the matroid (and thus $s=r$). As the hyperplane of our facet does not pass through $\mathbf{0}$, we can choose a normal vector $\ell$ such that $\ell\cdot\mathbf{a}_{i_{j}}=1$ for each $j=1,\dots,r$. Since $\ell\cdot\mathbf{0}=0$, this implies $\ell\cdot\mathbf{a}_{i}\leq 1$ for each $i\in E$.
Referring to the above, let us call a vector $\ell$ an admissible vector if $\ell\cdot\mathbf{a}_{i}\leq 1$ for each $i\in E$ and the elements for which equality holds form a full rank set in the matroid.
Proposition 4.1.
Let $A\in\mathbb{R}^{n\times|E|}$ be a totally unimodular matrix, representing a co-Eulerian regular oriented matroid $M$. Let $R$ be the linear span of the columns $(\mathbf{a}_{i})_{i\in E}$ of $A$. Then the extended root polytope satisfies
(4.1) $\tilde{\mathcal{Q}}_{A}=\left\{\,\mathbf{x}\in R\;\middle|\begin{array}[]{cl}h_{C^{*}}(\mathbf{x})\geq 0&\text{for all directed cocircuits }C^{*}\text{ of }G\\ \ell\cdot\mathbf{x}\leq 1&\text{for all admissible vectors $\ell$ of $G$}\end{array}\right\}.$
Here $h_{C^{*}}$ is the functional that appears in the definition of a cocircuit and we remark that our conventions guarantee that the restriction of $h_{C^{*}}$ to $R$ is determined by $C^{*}$. Similarly, even though the number of admissible vectors is typically infinite, there are only finitely many possibilities for the restriction of $\mathbf{x}\mapsto\ell\cdot\mathbf{x}$ to $R$.
Proof.
First note that each vector $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{A}$ also belongs to the right hand side. It suffices to check this for the generators of the convex hull: For any admissible vector $\ell$, by definition $\ell\cdot\mathbf{a}_{i}\leq 1$ for each $i\in E$, and $\ell\cdot\mathbf{0}=0\leq 1$, whence $\ell\cdot\mathbf{x}\leq 1$ for each $\mathbf{x}\in\tilde{\mathcal{Q}}_{A}$. Let $C^{*}$ be a directed cocircuit. Then by definition, the corresponding functional $h_{C^{*}}$ is such that $h_{C^{*}}(\mathbf{a}_{i})=0$ for $i\notin C^{*}$ and $h_{C^{*}}(\mathbf{a}_{i})=1>0$ for $i\in C^{*}$, moreover, $h_{C^{*}}(\mathbf{0})=0$. Hence each $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{A}$ satisfies $h_{C^{*}}(\mathbf{x})\geq 0$.
Conversely, we show that any vector $\mathbf{y}$, satisfying the conditions of the right hand side, belongs to ${\tilde{\mathcal{Q}}}_{A}$. Take any facet $F$ of ${\tilde{\mathcal{Q}}}_{A}$. If it does not contain $\mathbf{0}$, then by the argument before the statement of the proposition, $F$ lies in the hyperplane $\{\mathbf{a}\mid\ell\cdot\mathbf{x}=1\}$ for an admissible vector $\ell$, moreover, ${\tilde{\mathcal{Q}}}_{A}$ is a subset of $\{\mathbf{x}\in\mathbb{R}^{n}\mid\ell\cdot\mathbf{x}\leq 1\}$, wherefore the hyperplane of $F$ does not separate $\mathbf{y}$ from ${\tilde{\mathcal{Q}}}_{A}$.
If $F$ contains $\mathbf{0}$, then it must additionally contain a set of $r-1$ linearly independent vectors $S=\{\mathbf{a}_{i_{1}},\dots,\mathbf{a}_{i_{r-1}}\}$, where $r=\dim{\tilde{\mathcal{Q}}}_{A}={\rm rank}(M)$. Let $U\subseteq\{\mathbf{a}_{1},\dots,\mathbf{a}_{|E|}\}-S$ be the set of vectors that do not lie along $F$. Then $C^{*}=\{i\in E\mid\mathbf{a}_{i}\in U\}$ is a cocircuit, with $\{\mathbf{x}\mid h_{C^{*}}(\mathbf{x})=0\}$ being the linear span of $S$. As $F$ is a facet, all vectors $\mathbf{a}_{i}$ must lie on one side of the span of $S$. Thus $C^{*}$ must be a directed cocircuit, and we also see that $\{\mathbf{x}\mid h_{C^{*}}(\mathbf{x})=0\}$ supports $\tilde{\mathcal{Q}}_{A}$ along $F$. Again, we conclude that the hyperplane of $F$ does not separate $\mathbf{y}$ from ${\tilde{\mathcal{Q}}}_{A}$.
Altogether, if a vector $\mathbf{y}$ belongs to the right hand side of (4.1), then it is on the appropriate side of each facet-defining hyperplane, whence $\mathbf{y}\in{\tilde{\mathcal{Q}}}_{A}$. ∎
Next, we elucidate the connection of dijoins to the geometry of the extended root polytope. Analogously to the case of graphs, let us call a dijoin $K$ of a regular oriented matroid $M$ a long-arc dijoin if there is a signed circuit $C$ such that $|C^{-}|>|C^{+}|$ and $C^{-}\subseteq K$.
Proposition 4.2.
A point $\mathbf{p}\in\tilde{\mathcal{Q}}_{A}$ is in the relative interior of ${\tilde{\mathcal{Q}}}_{A}$ if and only if there exists a circuit-free dijoin $K$ of $M$ such that $\mathbf{p}=\sum_{i\in K}\lambda_{i}\mathbf{a}_{i}$, where $\lambda_{i}>0$ for each $i\in K$ and $\sum_{i\in K}\lambda_{i}<1$.
In the proof, we will use the Farkas lemma of totally unimodular matrices (see for example [8, Lemma 4.2.13]). Let us state the lemma in the form that we will need. For two matrices $P\in\mathbb{R}^{n\times a}$ and $Q\in\mathbb{R}^{n\times b}$, we denote by $(P|Q)\in\mathbb{R}^{n\times(a+b)}$ the matrix obtained by writing the columns of $Q$ after the columns of $P$.
Lemma 4.3 (Farkas lemma for TU matrices).
Suppose that $(P|Q)$ is a totally unimodular matrix, and $\mathbf{b}_{0}$, $\mathbf{b}_{1}$ are integer row vectors (of the appropriate dimension). Then there exists a row vector $\mathbf{x}$ such that $\mathbf{x}P=\mathbf{b}_{0}$, $\mathbf{x}Q\leq\mathbf{b}_{1}$ if and only if there do not exist column vectors $\mathbf{y}_{0}$ and $\mathbf{y}_{1}$ such that $P\mathbf{y}_{0}+Q\mathbf{y}_{1}=\mathbf{0}$, $\mathbf{b}_{0}\mathbf{y}_{0}+\mathbf{b}_{1}\mathbf{y}_{1}<0$ and $\mathbf{y}_{1}\geq\mathbf{0}$. Moreover, if there exist $\mathbf{y}_{0}$ and $\mathbf{y}_{1}$ satisfying the inequalities, then there also exist solutions $\mathbf{y}_{0}$ and $\mathbf{y}_{1}$ whose coordinates are in $\{0,1,-1\}$.
Proof of Proposition 4.2.
By Proposition 4.1, a point $\mathbf{p}\in{\tilde{\mathcal{Q}}}_{A}$ is in the interior of $\tilde{\mathcal{Q}}_{A}$ if and only if $f_{C^{*}}(\mathbf{p})>0$ for each directed cocircuit $C^{*}$, and $\ell\cdot\mathbf{p}<1$ for each admissible vector $\ell$. Recall that the functional induced by $C^{*}$ satisfies $f_{C^{*}}(\mathbf{a}_{i})=0$ whenever $i\notin C^{*}$, and $f_{C^{*}}(\mathbf{a}_{i})=1$ for each $i\in C^{*}$.
We start with the ‘if’ direction. Suppose that $K$ is a dijoin, and $\mathbf{p}=\sum_{i\in K}\lambda_{i}\mathbf{a}_{i}$ with $\sum_{i\in K}\lambda_{i}<1$ and $\lambda_{i}>0$ for each $i\in K$. Then $\mathbf{p}=\sum_{i\in K}\lambda_{i}\mathbf{a}_{i}+(1-\sum_{i\in K}\lambda_{i})\cdot\mathbf{0}$, in particular $\mathbf{p}\in{\tilde{\mathcal{Q}}}_{A}$. For any directed cocircuit $C^{*}$, we have $f_{C^{*}}(\mathbf{p})=\sum_{i\in C^{*}\cap K}\lambda_{i}>0$, since the intersection is nonempty (by the definition of a dijoin), and the summands are all positive. Similarly, for any admissible vector $\ell$ we have $\ell\cdot\mathbf{a}_{i}\leq 1$ for any $i\in E$, and $\ell\cdot\mathbf{0}=0$. Since the coefficient of $\mathbf{0}$ is positive, $\ell\cdot\mathbf{p}<1$. This implies that $\mathbf{p}$ is in the interior of ${\tilde{\mathcal{Q}}}_{A}$.
Now we prove the ‘only if’ direction. Let $\mathbf{p}$ be an arbitrary point in the relative interior of ${\tilde{\mathcal{Q}}}_{A}$, and take any convex combination $\mathbf{p}=\sum_{i\in S}\lambda_{i}\mathbf{a}_{i}+\mu\cdot\mathbf{0}$ with $\lambda_{i}>0$ for all $i\in S$, where $S\subseteq E$ is some subset. For a point of ${\tilde{\mathcal{Q}}}_{A}$, we can always find such a formula (usually more than one). We claim that if $\mathbf{p}$ is in the interior, then $S$ is necessarily a dijoin. Indeed, suppose that $S$ is disjoint from a directed cocircuit $C^{*}$. Then $f_{C^{*}}(\mathbf{p})=\sum_{i\in S}\lambda_{i}\cdot f_{C^{*}}(\mathbf{a}_{i})+\mu\cdot f_{C^{*}}(\mathbf{0})=0$, which would contradict $\mathbf{p}$ being an interior point of ${\tilde{\mathcal{Q}}}_{A}$.
Next, we show that either $\sum_{i\in S}\lambda_{i}<1$, or $S$ is a long-arc dijoin. Then, we show that if $S$ is a long-arc dijoin, then there exist another dijoin $S^{\prime}$ such that $\mathbf{p}=\sum_{i\in S^{\prime}}\lambda^{\prime}_{i}\mathbf{a}_{i}+\mu^{\prime}\cdot\mathbf{0}$ and $\sum_{i\in S^{\prime}}\lambda^{\prime}_{i}<1$. Lastly, we show that we can also make the dijoin circuit-free, which finishes the proof of the ‘only if’ direction.
If $\sum_{i\in S}\lambda_{i}<1$, then there is nothing to prove. So suppose that $\sum_{i\in S}\lambda_{i}=1$. We show that in this case $S$ is a long-arc dijoin.
The fact that $\mathbf{p}$ is an interior point of ${\tilde{\mathcal{Q}}}_{A}$ implies that there is no vector $\ell$ such that $\ell\cdot\mathbf{a}_{i}=1$ for each $i\in S$ and $\ell\cdot\mathbf{a}_{i}\leq 1$ for each $i\notin S$. Indeed, such a vector would give a supporting hyperplane of ${\tilde{\mathcal{Q}}}_{A}$ containing $\mathbf{p}$.
Take the totally unimodular representing matrix $A$ of $M$. Let us gather the columns corresponding to $S$ into the submatrix $A_{S}$, and the columns corresponding to $E-S$ into the submatrix $A_{E-S}$.
Our previous reasoning says that there is no row vector $\ell$ such that $\ell\cdot A_{S}=\mathbf{1}$, and $\ell\cdot A_{E-S}\leq\mathbf{1}$. (Here, we denote by $\mathbf{1}$ a row vector of appropriate dimension, with all coordinates equal to 1, and with a slight abuse of notation, we do not indicate the dimension).
As $A=(A_{S}|A_{E-S})$ is totally unimodular, by the Farkas lemma, if there exists no solution $\ell$ for the system of inequalities, then there exist $\mathbf{y}_{0}$, $\mathbf{y}_{1}$ such that $A_{S}\cdot\mathbf{y}_{0}+A_{E-S}\cdot\mathbf{y}_{1}=\mathbf{0}$, $\mathbf{1}\cdot\mathbf{y}_{0}+\mathbf{1}\cdot\mathbf{y}_{1}<0$, and $\mathbf{y}_{1}\geq 0$, moreover, $\mathbf{y}_{0}$ and $\mathbf{y}_{1}$ have all their coordinates in $\{0,1,-1\}$.
In other words, $\mathbf{y}=\begin{bmatrix}\mathbf{y}_{0}\\ \mathbf{y}_{1}\end{bmatrix}\in\ker(A)$. Recall that the circuits of $M$ are the minimal supports of elements of $\ker(A)$. In this case, there exists a set of signed circuits $\mathcal{C}_{0}$ with mutually disjoint supports such that $\mathbf{y}=\sum_{C\in\mathcal{C}_{0}}\chi_{C}$ (see for example [1, Lemma 4.1.1]). Hence $0>\mathbf{1}\cdot\mathbf{y}=\sum_{C\in\mathcal{C}_{0}}\mathbf{1}\cdot\chi_{C}=\sum_{C\in\mathcal{C}_{0}}(|C^{+}|-|C^{-}|)$. This implies that there is at least one circuit $C\in\mathcal{C}_{0}$ such that $|C^{-}|>|C^{+}|$. As $\mathbf{y}_{1}\geq\mathbf{0}$, elements of $(E-S)\cap C$ are in $C^{+}$. Hence $C^{-}\subseteq S$. Thus, $C$ witnesses the fact that $S$ is a long-arc dijoin.
Next, we show that if $\mathbf{p}=\sum_{i\in S}\lambda_{i}\mathbf{a}_{i}$ with $\lambda_{i}>0$, $\sum_{i\in S}\lambda_{i}=1$ where $S$ is a long-arc dijoin, then we can find a different dijoin $K$ such that $\mathbf{p}=\sum_{i\in K}\lambda_{i}\mathbf{a}_{i}$ with $\lambda_{i}>0$ and $\sum_{i\in K}\lambda_{i}<1$.
Let $C$ be the circuit that has $C^{-}\subseteq S$ and $|C^{-}|>|C^{+}|$. Such a circuit exists because of the definition of a long-arc dijoin. Let $\delta=\min\{\lambda_{i}\mid i\in C^{-}\}$ and let $S^{\prime}=\{i\in C^{-}\mid\lambda_{i}=\delta\}$. Then $\mathbf{p}=\sum_{i\in S-S^{\prime}\cup C^{+}}\lambda^{\prime}_{i}\mathbf{a}_{i}+\mu^{\prime}\cdot\mathbf{0}$ where
$\lambda^{\prime}_{i}=\left\{\begin{array}[]{cl}\lambda_{i}&\text{if $i\notin C$},\\ \lambda_{i}-\delta&\text{if $i\in C^{-}$},\\ \lambda_{i}+\delta&\text{if $i\in C^{+}$},\end{array}\right.$
and $\mu^{\prime}=\mu+(|C^{-}|-|C^{+}|)\cdot\delta$. This is a new convex combination for $\mathbf{p}$, where the coefficients are only positive for edges in $S-S^{\prime}\cup C^{+}$, and for $\mathbf{0}$. Moreover, $S-S^{\prime}\cup C^{+}$ is also a dijoin because any directed cut that intersects $C$ necessarily intersects both $C^{+}$ and $C^{-}$. Hence we have found a linear combination for $\mathbf{p}$ where now the coefficient of $\mathbf{0}$ is positive.
It remains to show that $S$ can be chosen circuit-free. If there is any circuit $C\subseteq S$, then orient it so that $|C^{-}|\geq|C^{+}|$. By the same argument as above, we can also write $\mathbf{p}$ as a combination of vectors of $S-S^{\prime}$, where $S^{\prime}$ is the set of elements of $C^{-}$ with minimal coefficients in the original linear combination. $S-S^{\prime}$ is still a dijoin, since $C^{+}\subseteq S-S^{\prime}$. Moreover, the coefficient of $\mathbf{0}$ can only increase by this operation. We can continue this until $S$ becomes independent in $M$, which happens in finitely many steps, as the size of the dijoin always decreases. ∎
Lemma 4.4.
Let $F$ be an independent set in the regular oriented matroid $M$, represented by the totally unimodular matrix $A$, and let $s\in\mathbb{Z}_{>0}$. A point $\mathbf{p}\in s\cdot{\tilde{\mathcal{Q}}}_{F}$ is a lattice point if and only if $\mathbf{p}=\sum_{i\in F}\mu_{i}\mathbf{a}_{i}$, where each $\mu_{i}$ is integer.
Proof.
This follows from Cramer’s rule and the total unimodularity of $A$. ∎
Proof of Theorem 1.6.
Let $K$ be a dijoin of cardinality $\nu(M)$. Then $\mathbf{p}=\sum_{k\in K}\mathbf{a}_{k}+\mathbf{0}$ is a point of $(\nu(M)+1)\cdot{\tilde{\mathcal{Q}}}_{A}$, moreover, it has integer coordinates. By Proposition 4.2, $\mathbf{q}=\frac{1}{\nu(M)+1}\mathbf{p}=\sum_{k\in K}\frac{1}{\nu(M)+1}\mathbf{a}_{k}$ is an interior point of ${\tilde{\mathcal{Q}}}_{A}$. Hence $\mathbf{p}$ is also an interior point of $(\nu(M)+1)\cdot{\tilde{\mathcal{Q}}}_{A}$.
We need to prove that for $s\leq\nu(M)$ there is no interior lattice point in $s\cdot{\tilde{\mathcal{Q}}}_{A}$. Suppose that $\mathbf{p}\in s\cdot{\tilde{\mathcal{Q}}}_{A}$ is an interior lattice point. Take $\mathbf{q}=\frac{1}{s}\mathbf{p}\in{\tilde{\mathcal{Q}}}_{A}$, which is then an interior point of ${\tilde{\mathcal{Q}}}_{A}$. By Proposition 4.2 there is a circuit-free dijoin $K$ such that $\mathbf{q}=\sum_{k\in K}\lambda_{k}\mathbf{a}_{k}$, $\lambda_{k}>0$ for each $k\in K$, and $\sum_{k\in K}\lambda_{k}<1$. In other words, $K$ is independent in $M$.
We can apply Lemma 4.4 for $s$, $K$, and $\mathbf{p}$. This tells us that for $\mathbf{p}$ to be an integer vector, $s\lambda_{k}$ needs to be an integer for each $k\in K$. Hence $s>\sum_{k\in K}s\lambda_{k}\geq|K|\geq\nu(M)$. ∎
Proposition 4.5.
The interior lattice points of $(\nu(M)+1)\cdot{\tilde{\mathcal{Q}}}_{A}$ are exactly the vectors that can be obtained as $\sum_{i\in K}\mathbf{a}_{i}$ for some minimal cardinality dijoin $K$.
Proof.
This statement can be proved the same way as Proposition 2.13. ∎
Proof of Theorem 1.7.
The statement follows from Theorems 1.6 and 2.12, and Proposition 4.5. ∎
One may wonder if an analogue of Theorem 3.1 holds.
Problem 4.6.
Given a regular matroid, which orientation has the interior polynomial of smallest degree? Is there an orientation whose interior polynomial is coefficientwise minimal?
5. Parking function enumerators and greedoid polynomials
In [20] it is proved that the parking function enumerator of an Eulerian digraph can be expressed as the interior polynomial of the cographic matroid. (Prior to that, [13] settled the planar case, which in turn extended [11, Corollary 5.9].) Hence the results of the previous section give us information on the degree of the parking function enumerator of an Eulerian digraph and, equivalently, on the degree of the lowest term of the greedoid polynomial. It turns out that similar results hold for all directed graphs.
In this section we recall the definition of a greedoid and the relationship between parking function enumerators and interior polynomials, then show how to generalize the result on the degree of the parking function enumerator to all directed graphs.
5.1. Preliminaries on greedoids and parking functions
Greedoids were introduced by Korte and Lovász as a structure in which the greedy algorithm works. Matroids are a special class of greedoids, but greedoids are able to express connectivity properties that matroids cannot.
Definition 5.1 (greedoid [14]).
A set system $\mathcal{F}$ on a finite ground set $E$ is called a greedoid if it satisfies the following axioms.
  • (1)
    $\emptyset\in\mathcal{F}$;
  • (2)
    for all $X\in\mathcal{F}-\{\emptyset\}$ there exists $x\in X$ such that $X-x\in\mathcal{F}$;
  • (3)
    if $X,Y\in\mathcal{F}$ and $|X|=|Y|+1$, then there exist an $x\in X-Y$ such that $Y\cup x\in\mathcal{F}$.
Elements of $\mathcal{F}$ are called accessible sets, and maximal accessible sets are called bases.
It follows from the axioms that bases have the same cardinality, which is called the rank of the greedoid.
An interesting subclass of greedoids is that of directed branching greedoids: For a digraph $G$ and its vertex $s$, the branching greedoid of $G$ rooted at $s$ is the set system consisting of the arborescences of $G$ rooted at $s$. The bases of this greedoid are the maximal arborescences.
$s$123469785
Figure 2. Eulerian digraph with root $s$. The non-dashed arcs form a spanning arborescence rooted at $s$.
The greedoid polynomial was introduced by Björner, Korte, and Lovász [5] in several equivalent ways. Here we recall the definition using activities with respect to a fixed ordering of the edges.
Let $B=\{b_{1},\dots,b_{r}\}$ be a basis of the greedoid. We can form words by concatenating the elements of $B$ in some order: $b_{i_{1}}b_{i_{2}}\dots b_{i_{r}}$. Such a word is called feasible if $\{b_{i_{1}},\dots,b_{i_{j}}\}\in\mathcal{F}$ for each $j=1,\dots,r$. Note that the axioms guarantee the existence of at least one feasible word for each basis. Let us fix an ordering of the ground set $E$. Now to any basis $B$ of the greedoid, one can associate its lexicographically minimal feasible word.
Example 5.2.
Consider the rooted digraph of Figure 2, and take the ordering of the edges indicated by the labelling. The edges $\{1,3,7,8,9\}$ form a spanning arborescence, i.e., a basis of the branching greedoid. The word $39187$ is feasible for this basis, but for example the word $89713$ is not (since $\{8\}$ is not an arborescence rooted at $s$). The lexicographically minimal feasible word is $13879$.
Definition 5.3 (external activity for greedoids [5]).
Let $(E,\mathcal{F})$ be a greedoid and fix an ordering of $E$. For a basis $B$, an element $e\notin B$ is externally active for $B$ if for any $f\in B$ such that $B\cup e-f\in\mathcal{F}$, the lexicographically minimal feasible word for $B$ is lexicographically smaller than the lexicographically minimal feasible word for $B-f\cup e$. The external activity of a basis $B$ is the number of externally active elements for $B$, and it is denoted by $e(B)$.
Definition 5.4 (greedoid polynomial, [5]).
Using the above, we associate
$\lambda(t)=\sum_{B:\text{ basis}}t^{e(B)}$
to an arbitrary greedoid.
We note that this is indeed well-defined, that is, independent of the ordering of the edges used to define the activities.
We will especially be interested in branching greedoids of digraphs. Swee Hong Chan proves [6] that the greedoid polynomial of a branching greedoid is a simple transformation of the enumerator of graph parking functions. Let us recall these notions, too.
Definition 5.5 (graph parking function).
For a directed graph $G$ and a fixed root vertex $s$, a graph parking function rooted at $s$ is a function $p\in\mathbb{Z}_{\geq 0}^{V-s}$ such that for each $S\subseteq V-s$, there is at least one vertex $u\in S$ with $p(u)<d(V-S,u)$, where $d(V-S,u)$ denotes the number of directed edges leading from $V-S$ to $u$.
We denote the set of these functions by ${\rm Park}(G,s)$. For a parking function $p\in{\rm Park}(G,s)$, we put $|p|=\sum_{v\in V-s}p(v)$.
Definition 5.6 (parking function enumerator).
For a directed graph $G=(V,E)$ and a fixed root vertex $s$, the parking function enumerator is the polynomial
${\rm park}_{G,s}(x)=\sum_{p\in{\rm Park}(G,s)}x^{|p|}.$
[…]
Figure 3. A rooted digraph (with root $s$), and its parking functions.
Example 5.7.
Figure 3 shows a rooted digraph and each one of its parking functions. Altogether, the parking funtion enumerator is $x^{2}+2x+1$.
The relationship of the greedoid polynomial and the parking function enumerator is the following.
Theorem 5.8.
[6, Theorem 1.3] $\lambda_{G,s}(x)=x^{|E|-|V|+1}{\rm park}_{G,s}(x^{-1})$
The second named author has previously observed the following connection.
Theorem 5.9.
[20] Let $G$ be a connected Eulerian digraph, and let $M$ be the directed dual matroid of $G$. Then
$\lambda_{G,s}(x)=x^{|E(G)|-|V(G)|+1}I_{M}(x^{-1})\quad\quad\text{and}\quad\quad{\rm park}_{G,s}(x)=I_{M}(x).$
Hence if $G$ is Eulerian, we can use Theorem 1.1 to obtain a formula for the degree of the parking function enumerator, or equivalently, a formula for the degree of the lowest term of the greedoid polynomial.
Proof of Theorem 1.8.
For the (directed) cographic matroid $M$ of $G$, a dijoin of $M$ (that is, a set of edges intersecting each directed cocircuit) corresponds to an edge set of $G$ that intersects each directed cycle. Hence dijoins of $M$ correspond to feedback arc sets of $G$. Now Theorem 5.9 implies the statement of the theorem. ∎
Remark 5.10.
The analogue of Theorem 5.9 is not true for general (non-Eulerian) digraphs. Indeed, for non-Eulerian digraphs, the parking function enumerator does depend on the root $r$, while the interior polynomial of the dual does not.
As an example, take the digraph $G$ that has two vertices $s$ and $u$, one edge from $s$ to $u$, and two edges from $u$ to $s$. As $G$ is planar, its dual is also a graphic matroid, associated to the digraph $G^{*}=(\{v_{1},v_{2},v_{3}\},\{\overrightarrow{v_{1}v_{2}},\overrightarrow{v_{2}v_{3}},\overrightarrow{v_{1}v_{3}}\})$. The unique parking function of $G$ rooted at $s$ associates 0 to $u$, whereby the parking function enumerator is ${\rm park}_{G,s}(x)=1$. However, it is easy to compute that $I_{G^{*}}(x)=1+x$.
Remark 5.11.
In [20], Theorem 5.9 is proved by noting that the complements of the arborescences of $G$ (that are bases in the cographic matroid $M$ of $G$) induce a triangulation of the root polytope of $M$. If we consider a non-Eulerian rooted digraph, then we can still associate the extended root polytope of $M|_{E-F}$ to any arborescence $F$, and this will be a simplex in ${\tilde{\mathcal{Q}}}_{M}$. If we take these simplices for each arborescence rooted at $s$, then they will be mutually disjoint, but they will typically not fill ${\tilde{\mathcal{Q}}}_{M}$. In fact, their union is not even necessarily convex. However it is easy to check that the union covers a neighborhood of $\mathbf{0}$ within ${\tilde{\mathcal{Q}}}_{M}$.
Take for example the planar digraph $G$ on two vertices $s$ and $u$, with two edges from $s$ to $u$ and three edges from $u$ to $s$. Then the dual can be given as $G^{*}=(\{v_{1},v_{2},v_{3},v_{4},v_{5}\},\{\overrightarrow{v_{1}v_{2}},\overrightarrow{v_{2}v_{3}},\overrightarrow{v_{3}v_{5}},\overrightarrow{v_{1}v_{4}},\overrightarrow{v_{4}v_{5}}\})$. In $G$ there are two arborescences rooted at $s$, and their respective complements in $G^{*}$ are $\{\overrightarrow{v_{1}v_{2}},\overrightarrow{v_{2}v_{3}},\overrightarrow{v_{3}v_{5}},\overrightarrow{v_{4}v_{5}}\}$ and $\{\overrightarrow{v_{1}v_{2}},\overrightarrow{v_{2}v_{3}},\overrightarrow{v_{3}v_{5}},\overrightarrow{v_{1}v_{4}}\}$. The resulting two simplices are the convex hulls of $\mathbf{0}$ and the vectors corresponding to these edges. Now the points $\mathbf{p}_{1}=\mathbf{1}_{v_{5}}-\mathbf{1}_{v_{4}}$ and $\mathbf{p}_{2}=\mathbf{1}_{v_{4}}-\mathbf{1}_{v_{1}}$ are in the respective simplices, on the other hand it is easy to check that $(\mathbf{p}_{1}+\mathbf{p}_{2})/2=(1/2)\cdot(\mathbf{1}_{v_{5}}-\mathbf{1}_{v_{1}})$ is not contained in either of the two simplices. That is, the union of these two simplices is not convex.
Nevertheless, Theorem 1.8 can still be generalized to any rooted digraph. For this, we first give a formula for the degree of the lowest term of the greedoid polynomial for general greedoids.
5.2. General greedoids
To give a formula for the degree of the lowest term of an arbitrary greedoid polynomial, let us make two easy observations.
Definition 5.12.
Let $X=(E,\mathcal{F})$ be a greedoid, and let $S\subseteq E$. We define $X|_{S}$ as the pair $(S,\mathcal{F}|_{S})$, where $\mathcal{F}|_{S}=\{A\in\mathcal{F}\mid A\subseteq S\}$, and call it the restriction of $X$ to $S$.
Claim 5.13.
$X|_{S}$ is a greedoid.
Proof.
It is easy to check that the axioms are true. ∎
Claim 5.14.
Let $X=(E,\mathcal{F})$ be a greedoid and $S\subseteq E$. Fix an ordering of the elements of $E$, and its restriction to $S$. Suppose that $F$ is a basis of both $X$ and $X|_{S}$. An element $e\in S-F$ is externally active for $F$ in $X|_{S}$ (with respect to the above mentioned ordering) if and only if $e$ is externally active for $F$ in $X$.
Proof.
The definition of external activity only considers bases that are subsets of $F\cup e$, and those are the same for $X|_{S}$ as for $X$. ∎
Theorem 1.12.
Let $X=\{E,\mathcal{F}\}$ be a greedoid of rank $r$. Let $k=\min\{|S|\mid S\subseteq E\text{ with }{\rm rank}(X|_{E-S})=r\text{ and }\lambda_{X|_{E-S}}(0)\neq 0\}$. Then in the greedoid polynomial of $X$, the coefficient of $x^{i}$ is zero for $i=0,\dots,k-1$ and the coefficient of $x^{k}$ is nonzero.
Proof.
Take an arbitrary basis $B$ of $X$, and let $P$ be the set of elements of $E-B$ that are externally passive for $B$. We claim that $B\cup P$ is a set such that ${\rm rank}(X|_{B\cup P})=r$ and $\lambda_{X|_{B\cup P}}(0)\neq 0$. The claim ${\rm rank}(X|_{B\cup P})=r$ follows immediately since $B$ is a basis of $X$. On the other hand, since elements of $P$ were all externally passive for $B$ in $X$, this remains so for $X|_{B\cup P}$, thus, in $X|_{B\cup P}$ there are no externally active elements for $B$. This shows that $\lambda_{X|_{B\cup P}}(0)\neq 0$. Hence $k\leq|E-B-P|$ for any basis $B$ of $X$. As the external activity of $B$ is $e_{X}(B)=|E-B-P|$, this shows that if the coefficient of $x^{i}$ is positive in the greedoid polynomial, then $i\geq k$.
It remains to show that the coefficient of $x^{k}$ is positive. Take a set $S\subseteq E$ with $|S|=k$ such that ${\rm rank}(X|_{E-S})=r$ and $\lambda_{X|_{E-S}}(0)\neq 0$. We show that there exists a basis $B$ such that $e_{X}(B)\leq|S|$. Since we have already proved that we cannot have $e_{X}(B)<k$, this will complete the proof.
As ${\rm rank}(X|_{E-S})=r$, each basis of $X|_{E-S}$ is a basis of $X$. Since $\lambda_{X|_{E-S}}(0)\neq 0$, the greedoid $X|_{E-S}$ has a basis $B$ such that $e_{X|_{E-S}}(B)=0$. That is, all elements of $E-S-B$ are externally passive in $B$. Hence, in $X$, only the elements of $S$ can be externally active for $B$, thus indeed, $e_{X}(B)\leq|S|=k$. ∎
5.3. Arbitrary directed graphs
In this section we use the result of the previous section to generalize Theorem 1.8 to arbitrary digraphs. Let $G$ be a digraph, and let $s$ be an arbitrary fixed vertex of $G$.
Note that when examining the greedoid polynomial, we can suppose that $G$ is $s$-root-connected, i.e., that each vertex of $G$ is reachable on a directed path from $s$. (This property is equivalent to $G$ having a spanning arborescence rooted at $s$.) Indeed, if $G$ is not $s$-root-connected, then the bases of the branching greedoid of $G$ rooted at $s$ will be the same as the bases of the branching greedoid of $G^{\prime}$ rooted at $s$, where $G^{\prime}$ is the subgraph of $G$ spanned by vertices reachable on a directed path from $s$. The edges of $G-G^{\prime}$ will be externally semi-active for any basis, and any edge ordering, that is in this case ${\rm park}_{G}(x)={\rm park}_{G^{\prime}}(x)$ and $\lambda_{G}(x)=x^{|E(G)|-|E(G^{\prime})|}\lambda_{G^{\prime}}(x)$. Hence from now on, we may suppose that $G$ is $s$-root-connected.
Swee Hong Chan points out that in the case of directed branching greedoids, the greedoid activity notion of [5] specializes to the following.
Definition 5.15 (external semi-activity in digraphs, [6]).
Let $G$ be an $s$-root-connected digraph with a fixed ordering of the edges. Let $A$ be a spanning arborescence rooted at $s$ in $G$. An arc $e\notin A$ is externally semi-active for $A$ if in the fundamental cycle $C(A,e)$ the maximal edge (with respect to the fixed ordering) stands parallel to $e$. If $e\notin A$ is not externally semi-active for $A$, then we call it externally semi-passive.
The name semi-activity comes from Li and Postnikov [15], who introduced this notion independently from the greedoid context.
Example 5.16.
For the rooted spanning arborescence in Figure 2, and the indicated edge ordering, edge number $2$ is externally semi-active. Indeed, its fundamental cycle is $\{2,3,8,7\}$, along which the maximal edge, $8$, stands parallel to $2$. On the other hand, $6$ is externally semi-passive, because its fundamental cycle is $\{6,1,3,8,7\}$, and among these, the maximal edge $8$ stands opposite to $6$.
Theorem 1.10.
Let $G=(V,E)$ be an $s$-root-connected digraph. The degree of the parking function enumerator of $G$ rooted at $s$ is equal to $|E|-|V|+1-{\rm minfas}(G,s)$.
Equivalently, for the greedoid polynomial of the branching greedoid of $G$ rooted at $s$, the coefficients of $x^{0},\dots,x^{k-1}$ are zero, and the coefficient of $x^{k}$ is nonzero for $k={\rm minfas}(G,s)$.
Recall that ${\rm minfas}(G,s)$ is the minimal cardinality of an edge set whose removal leaves an $s$-root-connected acyclic digraph (Definition 1.9). It follows from Theorems 1.8 and 1.10 that for a connected Eulerian digraph and an arbitrary vertex $s$, ${\rm minfas}(G)={\rm minfas}(G,s)$. It is also not very hard to prove this directly.
For general digraphs, ${\rm minfas}(G)$ and ${\rm minfas}(G,s)$ might differ. For example if $G$ has two vertices, $s$ and $v$, with one edge from $s$ to $v$, and two edges from $v$ to $s$, then ${\rm minfas}(G)=1$ but ${\rm minfas}(G,s)=2$.
Remark 5.17.
For the interior polynomial (consequently, also for the parking function enumerator of Eulerian digraphs), we had a definition using Ehrhart theory. For the parking function enumerator of general digraphs we are unaware of such a definition. Hence in the proof of Theorem 1.10, we need to use a different, combinatorial argument.
As we remarked in the introduction, Theorem 1.10 strengthens a theorem of Björner, Korte, and Lovász, who showed in [5] that for the branching greedoid of an $s$-root-connected digraph, the greedoid polynomial has a nonzero constant term if and only if the digraph is acyclic. They in fact prove more than this statement, also establishing the equivalence of this property to certain topological assumptions. For the sake of self-containedness, we give a short proof of the part of their statement that we will use.
Lemma 5.18.
[5] Let $G$ be an $s$-root-connected digraph. The greedoid polynomial $\lambda_{G,s}$ has nonzero constant term if and only if $G$ is acyclic.
Proof.
Suppose that $G$ is acyclic and fix an arbitrary ordering of its edges. In order to show that $\lambda_{G,s}(0)\neq 0$, we need to find a spanning arborescence (rooted at $s$) with 00 external activity.
Consider the spanning arborescence $A$ whose lexicographically minimal feasible word is lexicographically maximal. For any edge $e\notin A$, if there is any arborescence of the form $A^{\prime}=A\cup e-f$, then the lexicographically minimal feasible word for $A^{\prime}$ can only be smaller than that of $A$. Hence it suffices to show that there is always such an arborescence $A^{\prime}$. As $A$ is a spanning arborescence, it is a tree with a unique directed path from $s$ to any vertex. If we add an edge $e=\overrightarrow{uv}$ to $A$, then we create a unique cycle. Since we assumed that $G$ is acyclic, this cycle is not directed, in particular $v\neq s$. By removing the unique in-edge $f$ of $v$ that $A$ contains, we once again get a tree with exactly one directed path to each vertex, i.e., $A^{\prime}=A\cup e-f$ is an arborescence. With this, we have finished proving that $e(A)=0$.
Now we show that if the $s$-root-connected graph $G$ is not acyclic, then $\lambda_{G,s}(0)=0$. We do this by proving that for any spanning arborescence $A$, the externally semi-active edges form a feedback arc set of $G$, whereby if $G$ is not acyclic, the set of externally semi-active edges cannot be empty for any spanning arborescence.
Let $A$ be an arbitrary spanning arborescence (rooted at $s$), and let $C$ be an arbitrary directed cycle of $G$. We will work in the cycle space within $\mathbb{R}^{E}$, which is generated by the vectors $\sum_{e\in C^{+}}\mathbf{1}_{e}-\sum_{e\in C^{-}}\mathbf{1}_{e}$ taken for all cycles $C\subseteq G$. As the fundamental cycles of $A$ form a basis of the cycle space of $G$ over $\mathbb{R}$, the cycle $C$ can be written (with a slight abuse of notation) as a combination of some fundamental cycles: $C=\lambda_{1}\cdot C(A,e_{1})+\dots+\lambda_{j}\cdot C(A,e_{j})$, where each coefficient is nonzero. Some edges of these fundamental cycles might cancel out, but notice that $e_{1},\dots,e_{j}$ only occur in their respective fundamental cycles, wherefore they are all part of $C$. We may suppose without loss of generality that each $e_{i}$ is positive in $C(A,e_{i})$ and $C=C^{+}$. This implies $\lambda_{1}=\dots=\lambda_{j}=1$.
Let $e$ be the maximal edge in the union of the cycles $C(A,e_{1}),\dots,C(A,e_{j})$. If $e\in C$, then (since $C$ is a directed cycle) $e$ is parallel to $\{e_{1},\dots,e_{j}\}$ in $C$, and hence if $e\in C(A,e_{i})$, then $e$ is also parallel to $e_{i}$ in $C(A,e_{i})$, and it is the maximal edge in this cycle. Thus, $e_{i}\in C$ is externally semi-active for $A$. If $e\notin C$, then $e$ needs to occur in at least two fundamental cycles, once with positive, and once with negative sign. If $e$ occurs with positive sign in $C(A,e_{i})$, then its maximality in this cycle ensures that $e_{i}\in C$ is externally semi-active for $A$. In both cases we found an externally semi-active element in $C$, in particular the externally semi-active elements cover each directed cycle. ∎
Proof of Theorem 1.10.
We apply Theorem 1.12. If $G$ is $s$-root-connected, then the rank of the branching greedoid rooted at $s$ is equal to $|V(G)|-1$. For an edge set $S$, the rank of the branching greedoid of $G[E-S]$ remains $|V(G)|-1$ if and only if $G[E-S]$ is $s$-root-connected. On the other hand, Lemma 5.18 tells us that $\lambda_{G[E-S],s}(0)=0$ is equivalent to $G[E-S]$ being acyclic. Hence the condition of Theorem 1.12 indeed gives the condition of Theorem 1.10 for branching greedoids of root-connected digraphs. ∎